OpenStudy (anonymous):

Solve for x: quantity x minus 8 all over 3 equals 12

OpenStudy (anonymous):

(x - 8)/3 = 12 x - 8 = 3 * 12 x - 8 = 36 x = 24

OpenStudy (anonymous):

Ops, I mean x = 44

OpenStudy (anonymous):

Thank you soo much! :) @MelindaR

OpenStudy (anonymous):

Anytime!